OpenStudy (anonymous):

The half-life of a certain radioactive material is 38 days. An initial amount of the material has a mass of 497 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 4 days. Round your answer to the nearest thousandth. (1 point) A ; 462.029 kg B ; 0.087 kg C ; 1.040 kg D ; 0 kg

OpenStudy (wolf1728):

You'll find this page helpful: http://www.1728.org/halflife.htm Ending Amount = Beginning Amount / 2^n where n = number of half lives 4 days equals how many half lives (of 38 days)? 4/38 = 0.1052631579 half lives Ending Amount = 497 kg / 2 ^ 0.1052631579 Ending Amount = 497 kg / 1.0756905862 Ending Amount = 462.0287714392 kg Ending Amount = 462.029 (rounded) answer is A